OpenStudy (anonymous):

which of the points-slope equations below is correct for the line that passes through points (5,6) & (1,1)

jigglypuff314 (jigglypuff314):

point slope form: y - y1 = m(x - x1) when m = slope and given the point (x1, y1) so first find the slope, can you do that? :)

OpenStudy (anonymous):

sure okay, no problem

OpenStudy (anonymous):

y-1=5/4(x+)?

OpenStudy (anonymous):

I mean y-1=5/4(x+1) ?

jigglypuff314 (jigglypuff314):

that is correct :) Good Job! ^_^

OpenStudy (anonymous):

is there any more that I need to do? to find out you think

jigglypuff314 (jigglypuff314):

I believe y-1=5/4(x+1) is the answer

OpenStudy (anonymous):

see theres is more

jigglypuff314 (jigglypuff314):

ah right you can use the other point as well

OpenStudy (anonymous):

like which one?

OpenStudy (anonymous):

thanks so much

jigglypuff314 (jigglypuff314):

You're Welcome ^_^